Adv. Math. Econ. 8, 75-86 (2006)

Hedging bounded claims with bounded outcomes

Freddy Delbaen

ETH Zürich,Department of Mthematics,CH-8092 Zurich,Switzerland Received: January 31, 2005 Revised: October 10, 2005 JEL classification: G13, G11, C65 Mathematical Subject Classification (2000): 60G44, 91B16, 91B70 Abstract. We consider a financial market with two or more separate components each driven by a Brownian Motion. We look at the problem to hedge a bounded contingent claim in such a way that all the components remain bounded. The problem can also be rephrased as a problem in risk measures. Key words: hedging, infimal convolution, coherent utility functions, Fatou property, Brownian motion